# Text Indexing and Pattern Matching with Ephemeral Edits

### v1

@@ 2 Preliminaries
We consider finite strings over an integer alphabet $\Sigma=[0,\sigma)$. The elements of $\Sigma$ are called letters. For a string $T=T[0]\cdots T[n-1]$ over alphabet $\Sigma$, its length is $|T|=n$. For any $0\leq i\leq j<n$, the string $T[i]\cdots T[j]$ is called a substring of $T$. By $T[i\mathinner{.\,.}j]$ we denote its occurrence at (starting) position $i$, and we call it a fragment of $T$. When $i=0$, this fragment is called a prefix (we may denote it by $T[\mathinner{.\,.}j]$). When j=n1j=n-1, this fragment is called a suffix (we may denote it by $T[i\mathinner{.\,.}]$). We use $T[i\mathinner{.\,.}j)$ to denote the fragment $T[i\mathinner{.\,.}j-1]$ of $T$ or $T(i\mathinner{.\,.}j]$ to denote the fragment $T[i+1\mathinner{.\,.}j]$ of $T$.
@@
Given a set $\mathcal{S}$ of strings, a trie $\textsf{T}(\mathcal{S})$ is a rooted tree whose nodes represent the prefixes of strings in $\mathcal{S}$ [CHL07]. The prefix corresponding to node $u$ is denoted by $\textsf{str}(u)$; and the node $u$ is called the locus of $\textsf{str}(u)$. Sometimes we are interested in maximizing this prefix for an arbitrary string $S$. We say that the locus of an arbitrary string $S$ in $\textsf{T}(\mathcal{S})$ is the node $u$ such that $\textsf{str}(u)$ is a prefix of $S$ and $|\textsf{str}(u)|$ is maximized. The parent-child relationship in $\textsf{T}(\mathcal{S})$ is defined as follows: the root node is the locus of the empty string $\varepsilon$; and the parent $u$ of another node $v$ is the locus of $\textsf{str}(v)$ without the last letter. This letter is the label of the edge $(u,v)$. The order on $\Sigma$ induces an order on the edges outgoing from any node of $\textsf{T}(\mathcal{S})$. A node $u$ is branching if it has at least two children and terminal if $\textsf{str}(u)\in\mathcal{S}$.
A compacted trie is obtained from $\textsf{T}(\mathcal{S})$ by dissolving all nodes except the root, the branching nodes, and the terminal nodes. The dissolved nodes are called implicit while the preserved nodes are called explicit. The compacted trie takes $\mathcal{O}(|\mathcal{S}|)$ space provided that the edge labels are stored as pointers to fragments of strings in $\mathcal{S}$. Given the lexicographic order on $\mathcal{S}$ along with the lengths of the longest common prefixes between any two consecutive (in this order) elements of $\mathcal{S}$, one can easily compute the compacted trie in $\mathcal{O}(|\mathcal{S}|)$ time [KLA+01]. The suffix tree of a string $T$, denoted by $\textsf{ST}(T)$, is precisely the compacted trie of the set of all suffixes of $T$. Suffix trees usually come with suffix links: a suffix link $\textsf{slink}(u)$ for a branching node $u$ of $\textsf{ST}(T)$ is the node $v$ such that $\textsf{str}(v)$ is the longest proper suffix of $\textsf{str}(u)$, i.e., if $\textsf{str}(u)=T[i\mathinner{.\,.}j-1]$ then $\textsf{str}(v)=T[i+1\mathinner{.\,.}j-1]$. The suffix array of $T$, denoted by $\textsf{SA}(T)$, is the array of the suffixes of $T$ in lexicographic order. It is often useful to define the inverse suffix array as $\textsf{iSA}[\textsf{SA}[i]]=i$, for all $i\in[0,n)$. $\textsf{ST}(T)$, $\textsf{SA}(T)$, and $\textsf{iSA}(T)$ can be computed in $\mathcal{O}(n)$ time for any string $T$ of length $n$ over an alphabet $\Sigma=[0,n^{\mathcal{O}(1)}]$ [FFM00]. The edges of $\textsf{ST}(T)$ can be accessed in $\mathcal{O}(1)$ time if stored using perfect hashing [BCF+23].
@@
Let $T$ be a string of length $n$ over an integer alphabet $\Sigma=[0,\sigma)$. We would like to preprocess $T$ to support the following type of queries [GFB94]: for any $i,j\in[0,n)$ return all Occ occurrences of $T$ in $T[0\mathinner{.\,.}i]T[j\mathinner{.\,.}n)$. This type of query, which we denote here by $\textsf{PrefSuf}(i,j)$, can be simply and efficiently implemented due to the following recent result.
@@
For any string $T$ of length $n$ over an alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$, we can answer $\textsf{PrefSuf}(i,j)$ queries, for any $i,j\in[0,n)$, in $\mathcal{O}(1)$ time after an $\mathcal{O}(n/\log_{\sigma}n)$-time preprocessing. The data structure size is $\mathcal{O}(n/\log_{\sigma}n)$ and the output is given as a compact representation (a single arithmetic progression) of $\mathcal{O}(1)$ size.
In fact, the data structure of Theorem 4 [Pis25] can be constructed in $\mathcal{O}(n)$ time, for any alphabet $\Sigma$, if the longest common extension (LCE) data structure of [KK19] is replaced by the Z-algorithm [Gus97]. This is possible because one of the two indices for the longest common prefix queries asked is 0 and one of the two indices for longest common suffix queries is n1n-1.
@@
Assume we have a set $X$ of $n$ keys from a universe $U=[0,u)$ with a total order. In the predecessor problem, we are given a query element $q\in U$, and we are asked to find the maximum $p\in X$ such that $p\leq q$ (the predecessor of qq[NR21]. When $X$ is static, and its elements are given sorted, the $y$-fast trie data structure [Wil83] can be constructed in $|X|=\mathcal{O}(n)$ time and space, and it supports predecessor queries in $\mathcal{O}(\log\log u)$ time; see [NR21] for more details.
@@ 3 Text Indexing with Ephemeral Substring Edits
We design a data structure that supports pattern matching queries with ephemeral substring insertions, deletions, or substitutions. First, we preprocess the text $T=T[0\mathinner{.\,.}n)$ over alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$. Then, we can preprocess any arbitrary pattern $P=P[0\mathinner{.\,.}m)$ given online and allow any ephemeral sequence of edit operations on $T$. Before reverting the $i$th operation, we report all occurrences of $P$ in $T^{i}$ in $\mathcal{O}(\log\log n+\textsf{Occ})$ worst-case time. In Section 3.1, we present the preprocessing of the text, which occurs one time. In Section 3.2, we present the preprocessing of the pattern, which occurs one time per arriving pattern. In Section 3.3, we present the algorithm for processing the edit operations. Let us first start with describing the high-level idea.
@@
Suppose an edit operation $e_{i}$ occurs at some position of $T$. Our goal is to efficiently locate the longest prefix of $P$ ending right before the edit and the longest suffix of $P$ spanning the edit. It then suffices to trigger one prefix-suffix query [Pis25]. For locating the occurrences of $P$ not spanning the edit (they are before or after the edit, and so they are not affected by it), we use a standard approach from computational geometry. We focus on finding the longest such suffix of $P$ (finding the longest such prefix of $P$ is symmetric).
When a pattern $P$ arrives, we find the union of occurrences of the suffixes of $P$ in $T$ using the suffix tree of $T$. Although these occurrences can be $\Omega(m)$, we can encode them in at most $m$ intervals on $T$. These intervals are not pairwise disjoint because a suffix of $P$ can be a prefix of another suffix of $P$. We can, however, make them disjoint and still have $\mathcal{O}(m)$ pairwise disjoint intervals on $T$. The invariant is that every occurrence from $T$ is associated only with the longest suffix of $P$. Since the intervals are now disjoint, we can insert them in a predecessor data structure [NR21]. Hence, each edit takes $\mathcal{O}(\log\log n+\textsf{Occ})$ time.
@@ 3.1 Preprocessing the Text
The preprocessing of text $T$ is standard. We construct the suffix tree (with suffix links) $\textsf{ST}(T)$, the suffix array $\textsf{SA}(T)$, and the inverse suffix array $\textsf{iSA}(T)$. This takes $\mathcal{O}(n)$ time and space [FFM00]. We also store, at each node of $\textsf{ST}(T)$, the corresponding $\textsf{SA}(T)$ interval – this can be achieved in $\mathcal{O}(n)$ time using a bottom-up traversal on ST; inspect Fig. 1 for an example. We also construct a data structure to answer any range maximum query (RMQ) on SA and a data structure for answering any range minimum query (RmQ) on SA. This preprocessing takes $\mathcal{O}(n)$ time and space [BF00]. This concludes the preprocessing of $T$.
Figure 1: The suffix tree ST of $T=T[0\mathinner{.\,.}17)=\texttt{ananabannabanaana}$ with suffix links (dotted). The label of edges leading to leaf nodes is truncated after the first letter to avoid cluttering the figure. The suffix array SA of $T$ is [16,13,9,4,14,11,2,0,6,10,5,15,12,8,3,1,7][16,13,9,4,14,11,2,0,6,10,5,15,12,8,3,1,7] and is inferred from ST using an in-order DFS traversal. The node labeled 16, which represents suffix $T[16\mathinner{.\,.}17)$, stores also the SA interval $[0,8]$ and the node spelling string ban from the root stores the SA interval $[9,10]$.
@@ 3.2 Preprocessing the Pattern
We next sort the letters of $P$ in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space [Han04]. (Note that this sorting is required only when $\sigma=m^{\mathcal{O}(1)}$ does not hold.) Each letter of $P$ is then replaced by its rank; in particular, each letter is now an integer from $[0,m)$. We can thus construct the suffix tree $\textsf{ST}(P)$ of $P$ in $\mathcal{O}(m)$ time and space. Given $\textsf{ST}(P)$, we derive the following rooted tree, which we denote by $\textsf{TREE}(P)$; inspect Fig. 2 for an example. We dissolve every branching node $u$ in $\textsf{ST}(P)$ that does not represent a suffix of $P$ (along with its incoming edges) and connect every child $v$ of $u$ directly with the parent $\textsf{parent}(u)$ of $u$ via the edge $(\textsf{parent}(u),v)$ labeled by the concatenation of the labels of the edges $(\textsf{parent}(u),u)$ and $(u,v)$. Thus, the tree $\textsf{TREE}(P)$ consists of exactly $m+1$ nodes: the root node and one node per suffix of $P$ occurring in $T$. Every root-to-node path represents a single suffix. We have an internal (possibly branching) node if and only if it spells a suffix of $P$ that is a prefix of another suffix of $P$. The edge labels are encoded by intervals over $[0,m)$ and so the size of $\textsf{TREE}(P)$ is $\mathcal{O}(m)$ and it takes $\mathcal{O}(m)$ time to construct it using a traversal on $\textsf{ST}(P)$. This preprocessing takes $\mathcal{O}(m\log\log m)$ time (or $\mathcal{O}(m)$ time if we do not sort the letters of $P$) and $\mathcal{O}(m)$ space.
@@
Using the loci information obtained from $\textsf{ST}(T)$ (matching statistics), each node of $\textsf{TREE}(P)$ representing the suffix $P[i\mathinner{.\,.}m)$ is decorated with an interval $s,e]_{i}$. We then decompose the elements (indices) of node intervals (i.e., the union of occurrences of suffixes of $P$ in $T$) into a new collection of intervals such that $\textsf{iSA}[j]\in[s^{\prime},e^{\prime}]_{i}$ if and only if $P[i\mathinner{.\,.}m)$ is the longest suffix of $P$ occurring as a prefix of $T[j\mathinner{.\,.}n)$; inspect Fig. 2 for an example. This can be done for all node intervals using a traversal on $\textsf{TREE}(P)$. The total number of all new intervals obtained from $\textsf{TREE}(P)$ is still in $\mathcal{O}(m)$ because the number of new intervals is linear in the number of nodes and edges in the tree. By construction, these newly created intervals are pairwise disjoint. Moreover, by construction, these intervals are sorted because both $\textsf{ST}(T)$ and $\textsf{ST}(P)$ have been constructed using the same order: the one induced by the alphabet $\Sigma$. For every $s^{\prime},e^{\prime}]_{i}$ in the collection of intervals, we insert $s^{\prime}$ into a static predecessor data structure (a static $y$-fast trie [Wil83]) with satellite values $e^{\prime}$ and $i$. This preprocessing takes $\mathcal{O}(m)$ time and space because the $s^{\prime}$ values to be inserted are sorted.
@@
(b) $\textsf{TREE}(P)$ after making the intervals from the SA of $T$ disjoint.
@@
We also construct (at most) $|\Sigma|$ rooted trees, denoted by $\textsf{TREE}_{c}(P)$, for all $c\in\Sigma$ that occur in $P$; inspect Fig. 3 for an example. $\textsf{TREE}_{c}(P)$ is the tree induced by $\textsf{TREE}(P)$ such that every node in $\textsf{TREE}_{c}(P)$ represents a suffix whose preceding letter is $c$. All trees $\textsf{TREE}_{c}(P)$ can be constructed from the suffix tree $\textsf{ST}(P)$ of $P$ in $\mathcal{O}(m)$ total time. In one bottom-up traversal, we add the suffixes $P[i\mathinner{.\,.}m)$ with label c=P[i1]c=P[i-1] in a group labeled $c$. The suffixes in each group are added in lexicographic order. We then construct one compacted trie for each group of suffixes in linear time [KLA+01]. For each compacted trie, we dissolve the branching nodes that do not represent a suffix of $P$ (similar to $\textsf{TREE}(P)$) to obtain $\textsf{TREE}_{c}(P)$. We then decompose the node intervals in $\textsf{TREE}_{c}(P)$ into a new collection of intervals such that $\textsf{iSA}[j]\in[s^{\prime},e^{\prime}]_{i}$ if and only if $P[i\mathinner{.\,.}m)$ is preceded by $c$ and $P[i\mathinner{.\,.}m)$ is the longest suffix of $P$ occurring as a prefix of $T[j\mathinner{.\,.}n)$. This can be done in a way similar to $\textsf{TREE}(P)$ using a traversal on $\textsf{TREE}_{c}(P)$. The total size of all trees $\textsf{TREE}_{c}(P)$ is in $\mathcal{O}(m)$ because if a letter $c^{\prime}$ from $\Sigma$ does not occur in $P$ then no tree $\textsf{TREE}_{c^{\prime}}(P)$ is created. The total number of all intervals obtained from all trees $\textsf{TREE}_{c}(P)$ is in $\mathcal{O}(m)$. Due to how we construct them, the intervals from a single tree are pairwise disjoint and sorted. For every $s^{\prime},e^{\prime}]_{i}$ in the collection of intervals for $\textsf{TREE}_{c}(P)$, we insert $s^{\prime}$ into a separate static $y$-fast trie labeled by $c$ with satellite values $e^{\prime}$ and $i$. This preprocessing takes $\mathcal{O}(m)$ time and space.
@@
Let $\epsilon\geq 1$ be a constant given together with pattern $P$. It should be clear that the construction of $\textsf{TREE}_{c}(P)$ can be generalized to $\textsf{TREE}_{S}(P)$, for any string $S$ of length $|S|$. In particular, we construct the trees $\textsf{TREE}_{S}(P)$, for all $S$ with $|S|\in[1,\lfloor\epsilon\rfloor]$ in $\mathcal{O}(\epsilon m)=\mathcal{O}(m)$ time and space. This is possible because, for any fixed length $|S|$, there are $\mathcal{O}(m)$ suffixes of $P$ to be grouped; inspect Fig. 4 for an example.

### v2

@@ 2 Preliminaries
We consider finite strings over an integer alphabet $\Sigma=[0,\sigma)$. The elements of $\Sigma$ are called letters. For a string $T=T[0]\cdots T[n-1]$ over alphabet $\Sigma$, its length is $|T|=n$. For any $0\leq i\leq j<n$, the string $T[i]\cdots T[j]$ is called a substring of $T$. By $T[i\mathinner{.\,.}j]$ we denote its occurrence at (starting) position $i$, and we call it a fragment of $T$. When $i=0$, this fragment is called a prefix (we may denote it by $T[\mathinner{.\,.}j]$). When j=n1j=n-1, this fragment is called a suffix (we may denote it by $T[i\mathinner{.\,.}]$). We may use $T[i\mathinner{.\,.}j)$ to denote the fragment $T[i\mathinner{.\,.}j-1]$ of $T$ or $T(i\mathinner{.\,.}j]$ to denote the fragment $T[i+1\mathinner{.\,.}j]$ of $T$.
@@
Given a set $\mathcal{S}$ of strings, a trie $\textsf{T}(\mathcal{S})$ is a rooted tree whose nodes represent the prefixes of strings in $\mathcal{S}$ [CHL07]. The prefix corresponding to node $u$ is denoted by $\textsf{str}(u)$; and the node $u$ is called the locus of $\textsf{str}(u)$. Sometimes we are interested in maximizing this prefix for an arbitrary string $S$. We say that the locus of an arbitrary string $S$ in $\textsf{T}(\mathcal{S})$ is the node $u$ such that $\textsf{str}(u)$ is a prefix of $S$ and $|\textsf{str}(u)|$ is maximized. The parent-child relationship in $\textsf{T}(\mathcal{S})$ is defined as follows: the root node is the locus of the empty string $\varepsilon$; and the parent $u$ of another node $v$ is the locus of $\textsf{str}(v)$ without the last letter. This letter is the label of the edge $(u,v)$. The order on $\Sigma$ induces an order on the edges outgoing from any node of $\textsf{T}(\mathcal{S})$. A node $u$ is called branching if it has at least two children and terminal if $\textsf{str}(u)\in\mathcal{S}$.
A compacted trie is obtained from $\textsf{T}(\mathcal{S})$ by dissolving all nodes except the root, the branching nodes, and the terminal nodes. The dissolved nodes are called implicit while the preserved nodes are called explicit. The compacted trie takes $\mathcal{O}(|\mathcal{S}|)$ space provided that the edge labels are stored as pointers to fragments of strings in $\mathcal{S}$. Given the lexicographic order on $\mathcal{S}$ along with the lengths of the longest common prefixes between any two consecutive (in this order) elements of $\mathcal{S}$, one can compute the compacted trie in $\mathcal{O}(|\mathcal{S}|)$ time [KLA+01].
The suffix tree of a string $T$, denoted by $\textsf{ST}(T)$, is precisely the compacted trie of the set of all suffixes of $T$. Every terminal node in $\textsf{ST}(T)$ is labeled by the starting position of its corresponding suffix in $T$. Suffix trees usually come with suffix links by construction: a suffix link for a branching node $u$ of $\textsf{ST}(T)$, denoted by $\textsf{slink}(u)$, is the node $v$ such that $\textsf{str}(v)$ is the longest proper suffix of $\textsf{str}(u)$, i.e., if $\textsf{str}(u)=T[i\mathinner{.\,.}j]$ then $\textsf{str}(v)=T(i\mathinner{.\,.}j]$. The suffix array of $T$, denoted by $\textsf{SA}(T)$, is the array of the suffixes of $T$ in lexicographic order. It is often useful to define the inverse suffix array as $\textsf{iSA}[\textsf{SA}[i]]=i$, for all $i\in[0,n)$. The data structures $\textsf{ST}(T)$, $\textsf{SA}(T)$, and $\textsf{iSA}(T)$ can be constructed in $\mathcal{O}(n)$ time for any string $T$ of length $n$ over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$ [FFM00]. The edges of $\textsf{ST}(T)$ can be accessed in $\mathcal{O}(1)$ time if stored using perfect hashing [BCF+23].
@@
Let $T$ be a string of length $n$ over an integer alphabet $\Sigma=[0,\sigma)$. We would like to preprocess $T$ to support the following type of queries, known as prefix-suffix queries [GFB94]: for any $i,j\in[0,n)$ return all Occ occurrences of $T$ in $T[0\mathinner{.\,.}i]T[j\mathinner{.\,.}n)$. This type of query, which we denote here by $\textsf{PrefSuf}(i,j)$, can be simply and efficiently implemented due to the following recent result.
@@
For any string $T$ of length $n$ over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$, we can answer any $\textsf{PrefSuf}(i,j)$ query in $\mathcal{O}(1)$ time after an $\mathcal{O}(n/\log_{\sigma}n)$-time preprocessing. The data structure size is $\mathcal{O}(n/\log_{\sigma}n)$ and the output is given as a compact representation (a single arithmetic progression) of $\mathcal{O}(1)$ size.
In fact, the data structure of Theorem 4 [PIS25] can be constructed in $\mathcal{O}(n)$ time, for any alphabet $\Sigma$, if the longest common extension (LCE) data structure of [KK19] is replaced by the classical Z-algorithm [GUS97]. This is possible because one of the two indices for the longest common prefix queries asked is 0 and one of the two indices for longest common suffix queries is n1n-1.
@@
Assume we have a set $X$ of $n$ keys from a universe $U=[0,u)$ with a total order. In the predecessor problem, we are given a query element $q\in U$, and we are asked to find the maximum $p\in X$ such that $p\leq q$ (the predecessor of $q$) [NR21]. When $X$ is static, and its elements are given sorted, the $y$-fast trie data structure [WIL83] can be constructed in $|X|=\mathcal{O}(n)$ time and space, and it supports predecessor queries in $\mathcal{O}(\log\log u)$ time; see [NR21] for more details. Thus, when $u=n$, the query time becomes $\mathcal{O}(\log\log n)$.
@@ 3 Text Indexing with Ephemeral Substring Edits
In this section, we prove Theorem 1. We design a data structure that supports pattern matching queries with ephemeral substring insertions, deletions, or substitutions. First, we preprocess the text $T=T[0\mathinner{.\,.}n)$ over alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$. Then, we preprocess any arbitrary pattern $P=P[0\mathinner{.\,.}m)$ given online and support any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all occurrences of $P$ in $T^{i}$ in $\mathcal{O}(\log\log n+\textsf{Occ})$ worst-case time. In Section 3.1, we present the preprocessing of the text, which occurs one time. In Section 3.2, we present the preprocessing of the pattern, which occurs one time per arriving pattern. In Section 3.3, we present the algorithm for processing the edit operations. Let us first start with describing the high-level idea.
@@
Suppose an edit operation $e_{i}$ occurs at some position in $T$. Our goal is to efficiently locate the longest prefix of $P$ ending right before the edit and the longest suffix of $P$ spanning the edit. It is then sufficient to trigger one prefix-suffix query. For locating the occurrences of $P$ not spanning the edit (they are before or after the edit, and so they are not affected by it), we use a standard approach from computational geometry. We focus on finding the longest such suffix of $P$ (finding the longest such prefix of $P$ is symmetric).
When a pattern $P$ arrives, we find the union of occurrences of the suffixes of $P$ in $T$ using the suffix tree of $T$. Although these occurrences can be $\Omega(m)$, we can encode them using at most $m$ intervals on $T$. These intervals are not pairwise disjoint because a suffix of $P$ can be a prefix of another suffix of $P$. We can, however, make them disjoint and still have $\mathcal{O}(m)$ pairwise disjoint intervals on $T$. The invariant is that every occurrence from $T$ is associated only with the longest suffix of $P$. Since the intervals are now disjoint, we can insert them in a predecessor data structure. Hence, each edit takes $\mathcal{O}(\log\log n+\textsf{Occ})$ time.
@@ 3.1 Preprocessing the Text
The preprocessing of text $T$ is standard. We construct the suffix tree (with suffix links) $\textsf{ST}(T)$, the suffix array $\textsf{SA}(T)$, and the inverse suffix array $\textsf{iSA}(T)$. This takes $\mathcal{O}(n)$ time and space [FFM00]. We also store, at each node of $\textsf{ST}(T)$, the corresponding $\textsf{SA}(T)$ interval – this can be achieved in $\mathcal{O}(n)$ time using a bottom-up traversal on $\textsf{ST}(T)$; inspect Figure 1 for an example. We also construct a data structure to answer any range maximum query (RMQ) on $\textsf{SA}(T)$ and a data structure for answering any range minimum query (RmQ) on $\textsf{SA}(T)$. The preprocessing time and space is $\mathcal{O}(n)$ [BF00]. This concludes the preprocessing of $T$.
Figure 1: The suffix tree ST of $T=T[0\mathinner{.\,.}17)=\texttt{ananabannabanaana}$ with suffix links (dotted). The label of edges leading to leaf nodes is truncated after the first letter to avoid cluttering the figure. The suffix array SA of $T$ is [16,13,9,4,14,11,2,0,6,10,5,15,12,8,3,1,7][16,13,9,4,14,11,2,0,6,10,5,15,12,8,3,1,7] and is inferred from ST using an in-order bottom-up traversal. The node labeled 16, which represents suffix $T[16\mathinner{.\,.}17)$, stores also the SA interval $[0,8]$. The node spelling string ban from the root stores the SA interval $[9,10]$.
@@ 3.2 Preprocessing the Pattern
We next sort the letters of $P$ in $\mathcal{O}(m\log\log m)$ time and $\mathcal{O}(m)$ space [HAN04]. (Note that this sorting is required only when $\sigma$ is relatively large; i.e., $\sigma=m^{\mathcal{O}(1)}$ does not hold.) Each letter of $P$ is then replaced by its rank; in particular, each letter is now an integer from $[0,m)$. We can thus construct the suffix tree $\textsf{ST}(P)$ of $P$ in $\mathcal{O}(m)$ time and space. Given $\textsf{ST}(P)$, we derive the following rooted tree, which we denote by $\textsf{TREE}(P)$; inspect Figure 2 for an example. The tree $\textsf{TREE}(P)$ is on exactly $m+1$ nodes: the root node and one node per suffix of $P$. We start from $\textsf{ST}(P)$ and dissolve every branching node $u$ that does not represent a suffix of $P$ (along with its incoming edges) and connect every child $v$ of $u$ directly with the parent $\textsf{parent}(u)$ of $u$ via the edge $(\textsf{parent}(u),v)$ labeled by the concatenation of the labels of the edges $(\textsf{parent}(u),u)$ and $(u,v)$. Thus every root-to-node path represents a single suffix of $P$. We have an internal (possibly branching) node if and only if it spells a suffix of $P$ that is a prefix of another suffix (possibly many) of $P$. The edge labels are encoded by intervals over $[0,m)$ and so the size of $\textsf{TREE}(P)$ is $\mathcal{O}(m)$ and it takes $\mathcal{O}(m)$ time to construct it using a traversal on $\textsf{ST}(P)$. This preprocessing takes $\mathcal{O}(m\log\log m)$ time (or $\mathcal{O}(m)$ time if we do not sort the letters of $P$) and $\mathcal{O}(m)$ space.
(a) $\textsf{ST}(P)$; the nodes colored red are the ones which are going to be dissolved in $\textsf{TREE}(P)$.
(b) $\textsf{TREE}(P)$; the nodes colored red in $\textsf{ST}(P)$ are dissolved here.
Figure 2: $\textsf{ST}(P)$ and $\textsf{TREE}(P)$ for $P=P[0\mathinner{.\,.}11)=\texttt{mississippi}$. The label of edges leading to leaf nodes in $\textsf{ST}(P)$ is truncated after the first letter to avoid cluttering the figure. The labels are omitted in $\textsf{TREE}(P)$.
@@
Using the loci information obtained from $\textsf{ST}(T)$ (via the matching statistics algorithm), each node of $\textsf{TREE}(P)$ representing the suffix $P[i\mathinner{.\,.}m)$ is associated with the computed interval $s,e]_{i}$. Note that, by construction, the interval associated to every node is contained in the intervals associated to its ancestors. We then decompose the elements (indices) of node intervals (i.e., the union of occurrences of suffixes of $P$ in $T$) into a new collection of intervals such that $\textsf{iSA}[j]\in[s^{\prime},e^{\prime}]_{i}$ if and only if $P[i\mathinner{.\,.}m)$ is the longest suffix of $P$ occurring as a prefix of $T[j\mathinner{.\,.}n)$; inspect Figure 3 for an example. This can be done for all node intervals using a traversal on $\textsf{TREE}(P)$. The total number of all new intervals obtained from $\textsf{TREE}(P)$ is still in $\mathcal{O}(m)$ because the number of new intervals is linear in the number of nodes and edges in the tree. By construction, these newly created intervals are pairwise disjoint. Moreover, by construction, these intervals are sorted because both $\textsf{ST}(T)$ and $\textsf{ST}(P)$ have been constructed using the same order: the one induced by the same alphabet $\Sigma$. For every $s^{\prime},e^{\prime}]_{i}$ in the collection of intervals, we insert $s^{\prime}$ into a static predecessor data structure (in particular, a static $y$-fast trie [WIL83]) with satellite values $e^{\prime}$ and $i$. This preprocessing takes $\mathcal{O}(m)$ time and space because the $s^{\prime}$ values to be inserted are sorted.
@@
(b) $\textsf{TREE}(P)$ after making the intervals disjoint.
@@
We also construct (at most) $|\Sigma|$ rooted trees, denoted by $\textsf{TREE}_{c}(P)$, for all $c\in\Sigma$ that occur in $P$; inspect Figure 4 for an example. Specifically, $\textsf{TREE}_{c}(P)$ is the tree induced by $\textsf{TREE}(P)$ such that every node in $\textsf{TREE}_{c}(P)$ represents a suffix whose preceding letter is $c$. All trees $\textsf{TREE}_{c}(P)$ can be constructed from the suffix tree $\textsf{ST}(P)$ of $P$ in $\mathcal{O}(m)$ total time. In one bottom-up traversal, we add the suffixes $P[i\mathinner{.\,.}m)$, with c=P[i1]c=P[i-1], in a group labeled $c$. The suffixes in each group are added in lexicographic order. We then construct one compacted trie for each group of suffixes in linear time [KLA+01]. For each compacted trie, we dissolve the branching nodes that do not represent a suffix of $P$ (similar to $\textsf{TREE}(P)$) to obtain $\textsf{TREE}_{c}(P)$. We then decompose the node intervals in $\textsf{TREE}_{c}(P)$ into a new collection of intervals such that $\textsf{iSA}[j]\in[s^{\prime},e^{\prime}]_{i}$ if and only if $P[i\mathinner{.\,.}m)$ is preceded by $c$ and $P[i\mathinner{.\,.}m)$ is the longest suffix of $P$ occurring as a prefix of $T[j\mathinner{.\,.}n)$. This can be done in a way similar to $\textsf{TREE}(P)$ using a traversal on $\textsf{TREE}_{c}(P)$. The total size of all trees $\textsf{TREE}_{c}(P)$ is in $\mathcal{O}(m)$ because if a letter $c^{\prime}$ from $\Sigma$ does not occur in $P$ then no tree $\textsf{TREE}_{c^{\prime}}(P)$ is created. The total number of all intervals obtained from all trees $\textsf{TREE}_{c}(P)$ is in $\mathcal{O}(m)$. Due to how we construct them, the intervals from a single tree are pairwise disjoint and sorted. For every $s^{\prime},e^{\prime}]_{i}$ in the collection of intervals for $\textsf{TREE}_{c}(P)$, we insert $s^{\prime}$ into a separate static $y$-fast trie labeled by $c$ with satellite values $e^{\prime}$ and $i$. This preprocessing takes $\mathcal{O}(m)$ time and space.
@@
Let $\epsilon\geq 1$ be a constant given together with pattern $P$. It should be clear that the construction of $\textsf{TREE}_{c}(P)$ can be generalized to $\textsf{TREE}_{S}(P)$, for any string $S$ of length $|S|\geq 1$. In particular, we construct the trees $\textsf{TREE}_{S}(P)$, for all $S$ with $|S|\in[1,\lfloor\epsilon\rfloor]$ in $\mathcal{O}(\epsilon m)=\mathcal{O}(m)$ time and space. This is possible because, for any fixed length $|S|$, there are $\mathcal{O}(m)$ suffixes of $P$ to be grouped; inspect Figure 5 for an example.